# DeepVoid: A Deep Learning Void Detector

### v1

@@ 8 Conclusions
In this work, we present DeepVoid, a novel multi-class method of semantically segmenting the morphology of structure in cosmological simulations. We define four physically-motivated classes of large-scale structure based on the tidal tensor computed from the gravitational potential: voids, walls, filaments, and halos (Hahn et al., 2007). We then train deep convolutional neural networks to reproduce this truth table given a density contrast field as input. We train U-net models on the TNG300-3-Dark simulation. When trained on the full DM particle density field, with interparticle separation $\lambda=0.33$ $h^{-1}$ Mpc, the best model achieves a MCC of 0.81 and a void F1 score of 0.96 on validation data.
@@
While DeepVoid employs a physical definition of structure, some parameter choices still influence the segmentation results. The eigenvalue threshold for the tidal tensor has the most significant impact on structural classification. Setting it to zero restricts void classification to only the most underdense centers of voids, while larger thresholds expand void regions substantially, attenuating the connections between walls and filaments. We select a physically-motivated threshold that also matches previously observed volume fractions of voids.
@@
In the follow-up work to this paper, we will develop versions of DeepVoid that are trained on mock galaxy redshift surveys from simulations. These applications of DeepVoid will take as input survey format data and deal with all the accompanying complications, such as redshift-space distortions and handling survey mask boundaries. The redshift-space distortions can also be leveraged to improve training and therefore our predictions, since there is information inherent in the orientation of structures relative to the line-of-sight. Future iterations of DeepVoid will also incorporate galaxy properties that are known to vary with environment, such as color and luminosity, to examine how environmental dependence may improve void detection. Once DeepVoid models are trained on simulated survey data, the next step will be to apply them to galaxy survey data from SDSS (Abazajian et al., 2009b), SDSS BOSS (Dawson et al., 2013), and DESI (DESI Collaboration et al., 2016) to produce void catalogs suitable for cosmological tests.
9 Acknowledgements
Primary computing for this paper was done on Picotte, which is supported by Drexel’s University Research Computing Facility. Additional computing for this work was supported in part by the National Research Platform Nautilus cluster, which is supported in part by National Science Foundation (NSF) awards CNS-1730158, ACI-1540112, ACI-1541349, OAC-1826967, OAC-2112167, CNS-2100237, CNS-2120019, the University of California Office of the President, and the University of California San Diego’s California Institute for Telecommunications and Information Technology/Qualcomm Institute, a CONAHCyT Ciencia de Frontera grant CF-2023-1-1971 and UNAM PAPIIT grant IN115224. Thanks to CENIC for the 100Gbps networks.
@@ 9 Acknowledgements
Software: Tensorflow (Abadi et al., 2015), Keras (Chollet et al., 2015).
@@ Appendix A Tidal Tensor Derivation
To apply our eigenvalue-based segmentation scheme to DM particles in a simulated box, we first CIC the particle distribution to obtain the matter density contrast as described in Section 3. We can then relate this to the potential by use of the 3D Poisson equation:
@@ Appendix B Classification Performance Metrics
Most classification metrics are some combination of the elements of the confusion matrix, which we define in Section 6. To see examples of confusion matrices, see Figure 7.
Accuracy is one of the most commonly used and easily understood metrics for evaluating any predictive process. Specifically, accuracy is defined as the ratio of correctly classified instances to the total number of instances:
@@
where TP represents true positives, TN true negatives, FP false positives, and FN false negatives. In our case, we calculate the global accuracy of our classifier by macro-averaging, i.e. calculating the accuracy for each class and averaging the four scores. However, it is misleading to use accuracy as a metric for imbalanced datasets. In our case, where we have many more void voxels than filament or halo voxels (61% vs. 5% and 0.5% of the volume, respectively), the model could simply predict void for every voxel and reach what looks to be a reasonable level of accuracy, since false positives are not penalized.
@@
[…]
@@
An important metric for evaluating our models’ performance is the Sørenson-Dice coefficient or Dice score, which is also known in machine learning circles as the F1 score (Dice, 1945; Sørensen et al., 1948). The Dice score is the harmonic mean of precision and recall. For a binary classifier, it can be expressed as
@@
However, as this is a multiclass problem, we also want to evaluate our models’ prediction quality across all classes. There are several ways to average across classes: micro, macro, and weighted averaging. Micro-averaging calculates the F1 score globally by considering the total number of TP, FP, and FN across all classes. This approach is useful for dealing with imbalanced datasets, since it ensures that each instance contributes equally to the final F1 score. In contrast, macro-averaging simply consists of computing the F1 score for each class and averaging those scores, giving equal weight to each class. Weighted averaging is computed similarly to the macro-average, but weights each score by their ‘support’, or the number of true instances for each class. However, this may result in an F1 score that does not lie between the precision value and the recall value, obscuring the original meaning of the F1 metric.
One drawback of the F1 score is that TN are never considered in the calculation of the metric, and therefore it could be misleading for datasets with class imbalance such as ours. A metric that is often touted as a superior alternative is the Matthews correlation coefficient (MCC; Chicco & Jurman, 2020). The MCC was originally defined by Matthews (1975) but was reintroduced in a machine learning context by Baldi et al. (2000). The MCC provides a holistic view of the predictive performance of a classifier, even with imbalanced classes. It can be expressed by terms from the confusion matrix as:
@@
It ranges from $-1$ to $+1$, and can generally be thought of as a correlation coefficient with values at or near $-1$ meaning complete disagreement between prediction and ground-truth, 0 representing random guessing, while a score of $+1$ represents a perfect prediction. To receive a good MCC, a model must perform well on the positive and negative elements for each class.
Due to these considerations, the metrics that most accurately reflect the performance of DeepVoid models and that we use to differentiate between them are balanced accuracy, micro-averaged F1, void F1, and the MCC. In particular, the MCC succinctly captures the general quality of the segmentation across all classes, and the void F1 reflects how well we segment void regions, which are the primary concern of this work.
@@ Appendix C Batch normalization
An additional component we add to our U-nets that improves performance is batch normalization. First proposed by Ioffe & Szegedy (2015), batch normalization is meant to make training deep neural networks both faster and more stable by reducing the internal covariate shift. This term, while not strictly defined in the machine learning community, refers to the phenomenon where the distribution of inputs to each layer changes during training due to randomness in both the initialization of weights and inherent variability in the input data itself. Batch normalization aims to address this by normalizing the activations within each ‘mini-batch’ of the training set. It does this by computing the mean and the variance of the feature images, then scaling and shifting the normalized values using trainable parameters. If we take a mini-batch $\{x_{1},x_{2},...,x_{m}\}$ from a layer in the network, the mini-batch mean $\mu_{B}$ and the variance $\sigma^{2}_{B}$ are defined as:
@@
This normalization has been proven in a variety of applications to repress exploding gradients and to help prevent overfitting through its slight regularization effect. These effects have led to batch normalization to become a common tool in deep learning. Typically, the batch normalization layers are placed after the convolutional layer and before the activation function in each convolutional block. This placement ensures that the activations passed to subsequent layers are normalized. However, in our U-nets, we found that placing batch normalization layers in every block actually caused exploding gradients, harming classification quality and sometimes preventing the model from ever meaningfully learning. We experimented with other placements, and found that a batch normalization layer in every other block helped training significantly. Additionally, we tested other common network modifications such as L1 and L2 regularization and dropout layers, but neither impacted model performance positively.
@@ Appendix D Alternative Loss Function
While CCE is the most commonly used loss function for classification, there are several others that can be employed for multi-class classification. The focal loss, a form of CCE with two additional parameters $\alpha$ and $\gamma$, was created to address classification problems where one or more of the constituent classes is vastly over-represented (Lin et al., 2017). The focal loss takes the form
$L_{\mathrm{FOCAL}}(y,\hat{y})=-\sum_{i}^{N}\alpha_{i}\left(1-\hat{y_{i}}\right)^{\gamma}\log(\hat{y_{i}}),$ (D1)
where $y$ and $\hat{y}$ are the same as in the CCE function, $\alpha_{i}$ is a weight for class $i$, and $\gamma$ is a focusing parameter. The main difference between CCE and the focal loss is that CCE treats all classes equally and does not explicitly address class imbalance in the training dataset. The focal loss parameter $\gamma$ down-weights ‘easy’ examples, i.e. predictions that the model confidently classifies correctly, and focuses more on ‘hard’ examples. For datasets such as ours that have large class imbalances, the idea is that this will change the shape of the loss ‘landscape’ and ideally lead to improved performance.We tested several values of $\alpha$ and $\gamma$, starting with the default values of $\alpha=0.25$ (for all classes) and $\gamma=2$. Additionally, we set the weights of the $\alpha$ parameter to roughly reflect the balance of the classes in the datasets, so $\alpha=[0.6,0.25,0.15,0.1]$. Finally, we set the $\alpha$ weights to reflect which classes were most important to classify correctly, the void and wall classes. Therefore, we also set $\alpha=[0.6,0.5,0.25,0.25]$. For each of these cases, we also tested setting $\gamma=3$. While all of these models provide good segmentations, no base model with focal loss outperforms the SCCE versions.
We also examined using a combination of the CCE loss function and the Dice (F1) score (Galdran et al., 2022). CCE provides a strong baseline for voxel-by-voxel classifications that typically has stable gradients and easily extends to multi-label problems. The F1 score is much more robust to class imbalance and prioritizes correct shapes and boundaries. We found that models trained using a multi-class version of the F1 score often fail to get traction on the problem and do not train. Many weightings of these two loss functions are used in the literature. We tested using the average of the two. We found that using this combination loss resulted in base models that either matched or very slightly outperformed the SCCE models, but performed slightly worse than SCCE when using CL to train on sparser data.

### v2

@@ VIII Conclusions
In this work, we present DeepVoid, a novel multiclass method of semantically segmenting the morphology of structure in cosmological simulations. We define four classes of large-scale structures based on the tidal tensor computed from the gravitational potential: voids, walls, filaments, and halos (Hahn et al., 2007). We then train deep convolutional neural networks to reproduce this truth table, given a density contrast field as input. We train U-Net models on the TNG300-3-Dark simulation. When trained on the full DM particle density field, with interparticle separation $\lambda=0.33$ $h^{-1}$ Mpc, the best model achieves an MCC of 0.81 and a void F1 score of 0.96 on validation data.
@@
While DeepVoid employs a physical definition of structure, some parameter choices still influence the segmentation results. The eigenvalue threshold for the tidal tensor has the most significant impact on structural classification. Setting it to zero restricts void classification to only the most underdense centers of voids, while larger thresholds expand void regions substantially, attenuating the connections between walls and filaments. We select a physically motivated threshold that also matches previously observed volume fractions of voids.
@@
It is important to note that while DeepVoid was developed to detect cosmic voids in sparse datasets, it is also fully capable of identifying regions in walls, filaments and clusters (see Fig. 5). The confusion matrix (Fig. 7) shows the identification of regions in walls, filaments, and clusters even in sparse galaxy surveys. DeepVoid can enable environmental studies of galaxies in the cosmic web up to redshifts where other methods fail due to poor sampling of cosmic regions. The U-Net at the core of DeepVoid learns patterns in the density field and translates them to labels even for highly sparse samples. This is aided by the use of curriculum learning, in which the density field is progressively degraded, and the patterns are learned based on previous training.
In the follow-up work to this paper, we will develop versions of DeepVoid that are trained on mock galaxy redshift surveys from simulations. The identification of voids from biased tracers requires careful consideration, as the observed galaxy distribution can be affected by the background cosmology and (in the case of simulations used for training) the galaxy formation physics of the simulation or semianalytic model used to construct galaxy catalogs. These applications of DeepVoid will take as input survey-format data and deal with all the accompanying complications, such as redshift-space distortions and handling survey mask boundaries. The redshift-space distortions can also be leveraged to improve training and, therefore, our predictions, since there is information inherent in the orientation of structures relative to the line of sight. Future iterations of DeepVoid will also incorporate galaxy properties that are known to vary with environment, such as color and luminosity, to examine how environmental dependence may improve void detection. Once DeepVoid models are trained on simulated survey data, the next step will be to apply them to galaxy survey data from SDSS (Abazajian et al., 2009), SDSS BOSS (Dawson et al., 2013), and DESI (DESI Collaboration et al., 2016) to produce void catalogs suitable for cosmological tests.
Acknowledgements
We thank the anonymous referee for the suggestions, which substantially improved this paper. Primary computing for this paper was done on Picotte, which is supported by the Drexel University Research Computing Facility. Additional computing for this work was supported in part by the National Research Platform Nautilus cluster, which is supported in part by National Science Foundation (NSF) awards CNS-1730158, ACI-1540112, ACI-1541349, OAC-1826967, OAC-2112167, CNS-2100237, CNS-2120019, the University of California Office of the President, and the University of California San Diego’s California Institute for Telecommunications and Information Technology/Qualcomm Institute, a Consejo Nacional de Humanidades Ciencia y Tectnología (CONAHCyT) Ciencia de Frontera grant CF-2023-1-1971 and Universidad Nacional Autónoma de México (UNAM) Programa de Apoyo a Proyectos de Investigación e Innovación Tecnológica (PAPIIT) grant IN115224. Thanks to CENIC for the 100Gbps networks.
@@ Acknowledgements
Data availability
The code used in this paper is publicly available at the following link: https://github.com/skumagai3/deepvoid_misc. The data used to train and test DeepVoid can be provided upon reasonable request of the authors.
Software: Tensorflow (Abadi et al., 2015), Keras (Chollet and others, 2015).
@@ Appendix A Tidal Tensor Derivation
To apply our eigenvalue-based segmentation scheme to DM particles in a simulated box, we first CIC the particle distribution to obtain the matter density contrast as described in Section III. We can then relate this to the potential by use of the 3D Poisson equation:
@@ Appendix B Classification Performance Metrics
Most classification metrics are some combination of the elements of the confusion matrix, which we define in Section VI. To see examples of confusion matrices, see Figure 7.
Accuracy is one of the most commonly used and easily understood metrics for evaluating any predictive process. In particular, accuracy is defined as the ratio of correctly classified instances to the total number of instances:
@@
where TP represents true positives, TN true negatives, FP false positives, and FN false negatives. In our case, we calculate the global accuracy of our classifier by macroaveraging, i.e. calculating the accuracy for each class and averaging the four scores. However, it is misleading to use accuracy as a metric for imbalanced datasets. In our case, where we have many more void voxels than filament or halo voxels (61% vs. 5% and 0.5% of the volume, respectively), the model could simply predict void for every voxel and reach what looks to be a reasonable level of accuracy, since false positives are not penalized.
@@
$\text{balanced acc.}=\frac{1}{2}\left(\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}+\frac{\mathrm{TN}}{\mathrm{TN}+\mathrm{FP}}\right).$ (B2)
@@
An important metric for evaluating our models’ performance is the Sørenson-Dice coefficient or Dice score, which is also known in machine learning circles as the F1 score (Dice, 1945; Sørensen, 1948). The Dice score is the harmonic mean of precision and recall. For a binary classifier, it can be expressed as
@@
However, as this is a multiclass problem, we also want to evaluate our models’ prediction quality across all classes. There are several ways to average across classes: micro, macro, and weighted averaging. Microaveraging calculates the F1 score globally by considering the total number of TP, FP, and FN across all classes. This approach is useful for dealing with imbalanced datasets, since it ensures that each instance contributes equally to the final F1 score. In contrast, macroaveraging simply consists of computing the F1 score for each class and averaging those scores, giving equal weight to each class. Weighted averaging is computed similarly to the macroaverage, but weights each score by its “support,” or the number of true instances for each class. However, this may result in an F1 score that does not lie between the precision value and the recall value, obscuring the original meaning of the F1 metric.
One drawback of the F1 score is that TN are never considered in the calculation of the metric, and, therefore, it could be misleading for datasets with class imbalance such as ours. A metric that is often touted as a superior alternative is the Matthews correlation coefficient (MCC; Chicco and Jurman, 2020). The MCC was originally defined by Matthews (1975) but was reintroduced in a machine learning context by Baldi et al. (2000). The MCC provides a holistic view of the predictive performance of a classifier, even with imbalanced classes. It can be expressed by terms from the confusion matrix as:
@@
It ranges from 1-1 to +1+1, and can generally be thought of as a correlation coefficient with values at or near 1-1 meaning complete disagreement between prediction and ground truth, 0 representing random guessing, while a score of +1+1 represents a perfect prediction. To receive a good MCC, a model must perform well on the positive and negative elements for each class.
Due to these considerations, the metrics that most accurately reflect the performance of DeepVoid models and that we use to differentiate between them are balanced accuracy, microaveraged F1, void F1, and the MCC. In particular, the MCC succinctly captures the general quality of the segmentation across all classes, and the void F1 reflects how well we segment void regions, which are the primary concern of this work.
@@ Appendix C Batch normalization
An additional component we add to our U-Nets that improves performance is batch normalization. First proposed by Ioffe and Szegedy (2015), batch normalization is meant to make training deep neural networks both faster and more stable by reducing the internal covariate shift. This term, while not strictly defined in the machine learning community, refers to the phenomenon where the distribution of inputs to each layer changes during training due to randomness in both the initialization of weights and inherent variability in the input data itself. Batch normalization aims to address this by normalizing the activations within each “minibatch” of the training set. It does this by computing the mean and the variance of the feature images, then scaling and shifting the normalized values using trainable parameters. If we take a minibatch $\{x_{1},x_{2},...,x_{m}\}$ from a layer in the network, the minibatch mean $\mu_{B}$ and the variance $\sigma^{2}_{B}$ are defined as:
@@
This normalization has been proven in a variety of applications to repress exploding gradients and to help prevent overfitting through its slight regularization effect. These effects have led to batch normalization becoming a common tool in deep learning. Typically, the batch normalization layers are placed after the convolutional layer and before the activation function in each convolutional block. This placement ensures that the activations passed to subsequent layers are normalized. However, in our U-Nets, we found that placing batch normalization layers in every block actually caused exploding gradients, harming classification quality and sometimes preventing the model from ever meaningfully learning. We experimented with other placements and found that a batch normalization layer in every other block helped training significantly. Additionally, we tested other common network modifications such as L1 and L2 regularization and dropout layers, but neither impacted model performance positively.
@@ Appendix D Alternative Loss Function
While CCE is the most commonly used loss function for classification, there are several others that can be employed for multiclass classification. The focal loss, a form of CCE with two additional parameters $\alpha$ and $\gamma$, was created to address classification problems where one or more of the constituent classes is vastly overrepresented (Lin et al., 2017). The focal loss takes the form
[…]
where $y$ and $\hat{y}$ are the same as in the CCE function, $\alpha_{i}$ is a weight for class $i$, and $\gamma$ is a focusing parameter. The main difference between CCE and the focal loss is that CCE treats all classes equally and does not explicitly address class imbalance in the training dataset. The focal loss parameter $\gamma$ down-weights “easy” examples, i.e. predictions that the model confidently classifies correctly, and focuses more on “hard” examples. For datasets such as ours that have large class imbalances, the idea is that this will change the shape of the loss ‘landscape’ and ideally lead to improved performance. We tested several values of $\alpha$ and $\gamma$, starting with the default values of $\alpha=0.25$ (for all classes) and $\gamma=2$. Additionally, we set the weights of the $\alpha$ parameter to roughly reflect the balance of the classes in the datasets, so $\alpha=[0.6,0.25,0.15,0.1]$. Finally, we set the $\alpha$ weights to reflect which classes were most important to classify correctly, the void and wall classes. Therefore, we also set $\alpha=[0.6,0.5,0.25,0.25]$. For each of these cases, we also tested setting $\gamma=3$. While all of these models provide good segmentations, no base model with focal loss outperforms the SCCE versions.
We also examined using a combination of the CCE loss function and the Dice (F1) score (Galdran et al., 2022). CCE provides a strong baseline for voxel-by-voxel classifications that typically have stable gradients and easily extend to multi-label problems. The F1 score is much more robust to class imbalance and prioritizes correct shapes and boundaries. We found that models trained using a multiclass version of the F1 score often fail to get traction on the problem and do not train. Many weightings of these two loss functions are used in the literature. We tested using the average of the two. We found that using this combination loss resulted in base models that either matched or very slightly outperformed the SCCE models, but performed slightly worse than SCCE when using CL to train on sparser data.